From DIDEAS Wiki
- 1 Oct 2009
- 1.1 todo
- 1.2 sat oct 31
- 1.3 notes
- 1.4 fri oct 30
- 1.5 thur oct 29
- 1.6 wed oct 28
- 1.7 tue oct 27
- 1.8 mon oct 26
- 1.9 sun oct 25
- 1.10 sat oct 24
- 1.11 fri oct 23
- 1.12 thur oct 22
- 1.13 wed oct 21
- 1.14 tue oct 20
- 1.15 mon oct 19
- 1.16 sun oct 18
- 1.17 notes
- 1.18 sat oct 17
- 1.19 fri oct 16
- 1.20 thu oct 15
- 1.21 wed oct 14
- 1.22 tue oct 13
- 1.23 mon oct 12
- 1.24 sun oct 11
- 1.25 sat oct 10
- 1.26 fri oct 9
- 1.27 thr oct 8
- 1.28 wed oct 7
- 1.29 tue oct 6
- 1.30 mon oct 5
- 1.31 sun oct 4
- 1.32 sat oct 3
- 1.33 Fri Oct 2
- 1.34 Thur Oct 1
- 1.35 notes
- order 10k arrays from digikey.!!!!!!!!
- measure drop across D14 (pg 11) SC/MC
- function generator for Sat
- get LF potential function
- for LF, if ABS - INC encoder error, then don't drop into mode 0!
- add "text code" the ASC output from SWIFI
- review schematics for MPD
sat oct 31
- further improved keys.py, documenting. PF_EEPROM#keys.py
- develop bench_prep.py for EEPROM config.
fri oct 30
- 11:30-2pm -15
- review schematics, assist GG w/reprogram of SWIFI
- 3:30p - 7:25p -1hr
- assist GG w/matlab
- testing out ICD3 programmer
- develop key-value system for EEPROM
- 10p - 2:20a -45
- write python for key-value, and test.
thur oct 29
- 12pm - 3pm -15min
- review IMU schematics, make changes for low power
- work w/CA to get bench testing running
- AG tuning., +20 phone to MPD, 45 min BT to AG robot
- discuss status w/RC
- improve the benchtest program, and SC app
- start overnight stress test w/fan
- reprogram SWIFI for lab. collect boards for fri 3pm tuning.
wed oct 28
- 1 hr home
- work w/CA to debug boards
- review IMU schematics, discuss w/ZH BW requirements.
- meet w/GG
tue oct 27
- 3:30pm - 5:05
- explain to GG and CA programming of boards
- 6:30-3:35 -30 mins
- ordered realice and ICD3
- assist CA w/debug
- work on testjig tags.
- write benchtest.py
mon oct 26
- 20 mins - emails to MR re Rev 300.
- 6pm - 3:40am (-1 hr dinner w/rc)
- explain MC operation to CA
- discussion w/GG
- discussion w/RC
- review PFF history, notice 13hz BW on torque for present PFF
- added boot retry count, that stops after 5 attempts.
- start to develop sensor to detect SG failure, the naive HP filter won't work...
sun oct 25
- 30 mins - looking for printf source, ftoa source, designing async emu
- complete integration of printf and async emulation into SWFI
- tested async emulation at 38400 k buad, and 19xxx - TX worked at the 38.4k, but Rx only worked at upto 19.xk
sat oct 24
- 1:30p-2:40am -15
- discuss w/RC project status, moving forward. Needs list for CA
- discuss w/RC MPD schematics, get IMU DS 
- reply to GG email
- start architecture docs.Pf_users
- test baudrates on 2.6.31 kernel - find problems send email to ED
- locate and test a new SWIFI
- wire S3 for development
- measure RTC clock coin cell voltage on two boards, create ECO #21
- SWIFI 4c operational, discovered that vert 1mm headers are backwards, replaced.
- experiment with async emulation @ 9600 baud
- experiment w/S3. Issues. having encoder problems, could be new firmware.
fri oct 23
- 12:40 - 1:30
- schematic review
- conference call re S03
- 4:35p - 9:32pm -30 mins
- schematic review, seemed to have made it though this. made note on the ECO pages.
- look into eddy current sensors and alternative magnetic sensors to replace linear hall
thur oct 22
- 1:00pm -2:10
- generate correct camera artwork for wip4c1, upload to SVN
- 3:00pm - 8:45p
- pick up components from lab
- prepare for tuning trial
- turning trial, review data; find there was a SG offset thoughout. caused a false LS trigger during swing.
- work w/ED on solidworks and CA on getting router ready, etc.
- 12:00 - 2am
- work on MPD schematics.
wed oct 21
- 1pm - 6:30; - 1.5 hrs
- working on invoice
- emails / phone discussions.
- discuss re bat and power infrastructure / PB switch;
- 5:40 - 6:30, discuss firmware status and track study ~10 days; 219 results and SWIFI packet loss; email to document firmware issue, review changes on SVN
- create docs and BOM and order list for dclamp2
- re-create docs for swifi wip4a.
- review MPD scheamtics.
tue oct 20
- 12:30p - 6:30p -45
- discuss w/HH, DG re firmware
- write software to locate serial port for lifefix.
- 10:30p - 3:10a
- fixed lifefix process list issue
- energy calculations for lifefix
mon oct 19
- 10:45a-12:30p phone calls from home w/ED and life fix.
- arrive 1:15-4:00
- meeting for 1.30 hr with DG and ED
- meet w/ZH - discuss time data in datalogs.
- replace signal handle module and get LF running again.
- document bugs in MC, looking for cause of qV overflow.
- work on abs encoder bug fix.
- more conditional LF compilation
- try to real-time plot LF log.
sun oct 18
- 2:00 pm - 4:00 pm
- fixing baudrate problems in call code.
- consider improvements to the EEPROM init table.
- 4:40- 2:50am - 30 mins
- emailed MR re status of D13.
- emailed DG re computer setup, ED re DNS for SWIFI
- python API for reading sensors.
- tested Pser2h #1, removed EEPROM, tried ISP selection with DTR - will be difficult to make use of
- got LF running
sat oct 17
- 50 mins : documenting / coding new state machine
- 2:15p - 9:00p
- meet w/DG and RC - discuss LF status and need for code, 1pm meeting Monday w/ED
- work w/CA till 7:30. We go over testing of boards, measurement of D13
- setup of computer for CA, reinstall my account. Create wiki account for him
- setup digital scope to verify baud rate.
- corrected baud rate code in wip4a
- disabled auto reset WIPORT feature in run_lifefix
- reprogrammed SWIFI on LF w/corrected code. started mods to MC.
- Pomona Electrionics - Model 6341 - very nice voltage probes.
- FT2232H DS
fri oct 16
- 1?? hr : flushing out details for new statemachine. (improvements to SW / ES w/goal of estimating speed)
- developing W statemachine ideas that will allow for softer K3 at speed - and to enter LS after FF
- 45 mins : create SWIFI assembly documentation, explain assembly with CA
- create assembly docs for Wip4c
thu oct 15
- email re LF crash, working w/CA, ED, and DG on various LF issues, getting it going, etc.
- 3:50p - 1:30a
- power computations for DG re TB hall trial. 42s, 56.5J to MOT, 19.6J=IIR, 7.3=BAT ESR
- w/ED review FTDI driver source to see about faster baud rates
- test high baud rates w/do_command. seems to support upto 3Mbaud w/DC - however dataport has many many errors, and appears to be dropping data even with a simple high speed reader.
- very confused by datarate of MC. 49 bytes in 600uS, 8 bits in 9.9uS! - but should be 1Mbaud.
- both the SWIFI and the MC/SC all have messed up high speed baudrates.
- cu = 0.0039; temperature = (res - r0)/(r0 * cu);
- sent plot to RC from the OCt13 trail
- created tags entry for the OCt13 "SPLT3" tuning.
- documentation for Pser2h and ldwash2
wed oct 14
- 1:20pm - 7:50p
- work with GA on IMU diagnoses, he finds cold solder join
- discuss data set and review trial w/ZH
- discuss documentation w/DG and DS
- discuss power consumption with Dk
- discuss PFF w/DG, then review LF source of oscillation. Adding compliance should reduce the problem
- modifiy LF to stop on ESC, modify LF to not halt on matlab requirement. (untested)
tue oct 13
- 11:20- 12:20
- locate computer for CA
- write T style log class for lifefix.
- arrive 1:05p - 11:30pm
- tuning for subject.
- move lifefix
- place digikey order for 2nd lifefix components.
mon oct 12
- 2:10p - 12:30a
- matlab running w/IMU dataset, added cmd line option to wifi_fast to select .vlist file
- verified PP side of FF, cmd 1445 w/1 to enable, cmd 1444 reads back; present values in MC are far too aggressive for PP
- unable to get load/save defaults to work from keyboard UI, cmd 146,147 work.
- need (2) SC-SWIFI cables, no spares find one, diagnose that has PWR-GND reverse on 6pin SWIFI side.
- MC (S02)not able to complete commutation
- tuning on Hugh, explain operation to DG
- improvements to EEPROM init table
- built calb tables from ZH email
- loaded calb tables into S1
- sg_tor_const needs to be set by calb table
- stop by lab, get PCBs, and trim.
- cut PCBs, respond emails
- review MPD R200, make notes, RC wants Red PCBs
sun oct 11
- 3pm - 7pm, 1.5 hrs re tools
- prepare for, and restore after for power down
- working on LF, and run out of office by bld manager. sort tools in the interim
- get lifefix operational
- 9:30- 2:20a
- review w/EP how to improve graphic output speed for LF
- rewire S2 for development.
- MC code conditionally compiles
- robo_config now returns error notices about invalid tables from AK
sat oct 10
- 2 hrs
- attempting access to LF and pedant. Emails to eric
- status and task list for tech.
- discuss status w/RC
- repair Windows box config - fix dual monitor - disable desktop search
- assess status of SP robots, neither is configured correctly.
- working on lifefix users manual : Lifefix_users
- start up lifefix
fri oct 9
- 12:15 - 1:20pm
- review LF data, compose possible hypothis to cause of problem
- develope parts lists/BOM for orders
- 5-9:05 -20
- try to recover LF datafiles from linux box - not able to get samba or SSH access
- develope parts lists/BOM for orders
- place digikey order
- contact PCB fab re shipping to new location
thr oct 8
- 2:00p - 11:30
- help ED to run LF, UDP term needs to be running
- work out how UPS will interface to system. power off: 7pm-11pm Fri, 5pm-?? Sun
- investigating tracking problems with LF.
- getting LF running, create several new envelopes
- make improvements (fix bugs) to run_lifefix.
wed oct 7
- develop parts list for latest PCB fab project
- 6:15 - 8:40
- explain to DG, test, correct backup ankle firmware. SWIFI has problem sending garbage
- test PDA, find working BT. Need to locate 2 others.
- charge batteries
- 11:05 - 2:45
- test WOW ankles, replace bullet connector (test others), verify PDA operation
- run LF, find problems with envelop.
- WIP4C - change 8x3.55 cap to 220uF
- review PN for bussed resistors
- find a sub for the value 5.87k (5.90k?)
- find value for QY1
- change 22uh to 22u
tue oct 6
- install, experiment with eispice for simulation
- 2:40p - 11:30 -25
- discuss project status w/ZH
- meet w/DG - explain VS and KS loops, LF, we study backlash and photo disassembly
- set up LF photos on webserver, setup software files for GG
- investigate SG problem on TJ
- review SG layout
mon oct 5
- 30 mins - developing switching power simulation
- 3:15- 5:45, -15
- review IW0905 DRC errors, correct and submit
- PCB submission accepted from SPE
- ordered storage bins
- 6:15pm - 10:45, -45 mins
- review time drift experiment, improve translation of log file to matlab.
- meet w/RC re new hires, plans for each, etc
- review MPD hardware design changes, email MR
sun oct 4
- 30 mins - developing switching power simulation
- 4:30pm- 7:20pm
sat oct 3
- 1pm - 2:45
- analysis of LF log file (update program for new log format, correct errors.)
- 3:55 - 8:45pm (estimate but last email at 7.pm)
- LF stopped after 11014 passes - review physical status of LF, appears okay.
- the most direct observation of the cause of stop is that cycle timer counter stopped after 500 tics. Could be that SWIFI data wasn't received yet, or that the timer actually stopped!
- belt is slipping and which causes the zero to drift - which causes the encoder trigger point to be missed.
Fri Oct 2
- 11:30a - 1:15p
- review AVR report from PCB fab house. Errors need to be correct and resubmitted.
- 5:00p - 7:00 p
- correct DRC mistakes, resubmit.
- LF stopped after 314278 passes (22:14pm)
- 1am - 2am??
- review LF stopped data, and restart
Thur Oct 1
- improve gnd grid, add plane for switcher, and correct poor routing issues w/WIP4C1.
- improve manual routing of WIP4C1, manual route of VCC33, re run auto route
- create a dclamp2 -
- submit project to fab house for DRC analysis
- 9:30p - 2:15 a
- review time drif experiment and improve output and start plotting
- test LF, checking output data, build new envelope for 2000N
- setup testjig to monitor copley (required reboot and serial cable was moved)
- running lifefix w/trajectory enevelope testing and SWIFI logging to linux box.